# Elimination distance to bounded degree on planar graphs

### v3

@@ 3 The proof
Then, assume that $b$ is in $A_{\omega}\cup\bigcup\limits_{i\leqslant\omega}T_{i}$. Then $b$ is both $\leqslant_{G}$ and $\leqslant_{H}$-maximal. Additionally, $S_{b}^{G}=S_{b}^{H}$ if $b$ is not adjacent to $a$, and $S_{b}^{G}=S_{b}^{H}\cup\{a\}$ otherwise. Note that as $a$ is $(k,d)$-safe, it has no neighbor of degree $d+1$, hence, if $b$ is adjacent to $a$ it has degree at most $d-1$ in $H$, and $|S_{b}^{H}|\leqslant d-1$ so $|S_{b}^{G}|\leqslant d$. For the last point of Definition 2, note that for any $v\in S^{G}_{b}$, $v$ also belongs to $A_{\omega}\cup\bigcup\limits_{i\leqslant\omega}T_{i}$, we then have $\{w~{}:~{}w<_{G}v\}=\{w~{}:~{}w<_{G}b\}=(m_{i})_{i<\omega}$.
Finally, we look at the case where $b$ is not in $A_{\omega}$, is not $m_{i}$ nor in $T_{i}$ for any $i\leqslant\omega$. In this case, we have that $b$ cannot have neighbors in $A_{\omega}$ nor any of the $T_{i}$ for $i\leqslant\omega$. To see this, assume that there is a $v\in T_{i}$ neighbor to $b$. This implies that $b$ is in $A_{i-1}$, as it is connected to $v$, the latter being in $A_{i-1}$, which is the connected component of $a$ in $G-\{m_{1},\ldots,m_{i-1}\}$. As $b\not\in T_{i}$, we have $m_{i}\leqslant_{H}b$ and $m_{i}\not\leqslant_{H}v$ which contradict that $\leqslant_{H}$ is an elimination order. This contradiction also holds if $b$ has a neighbor in $A_{\omega}$ as this would imply that $b\in A_{\omega}$.
@@
To conclude, we have that $\leqslant_{G}$ is indeed an elimination order to degree $d$ of depth $k$ for $G$. This ends the proof that $a$ is irrelevant. ∎
We can now prove our main result, Theorem 1.
Proof of Theorem 1.
Let $k,d$ be two integers and $G$ be a connected $K_{5}$-minor free graph. We set $h(k):=(4k+7)\cdot k$. Let $g(\cdot)$ be the function from Theorem 5.
First run the algorithm of Theorem 5 with parameter $h(k)$ (or, if $G$ is planar, use the algorithm of Theorem 6).
@@
The algorithm outputs a tree-decomposition of width $g(h(k))$, or of width $5h(k)$ if $G$ is planar. We then use Courcelle’s Theorem (Theorem 7) to decide whether $G\in\mathscr{C}_{k,d}$.
@@
If more than $k^{2}$ branch sets contain a node of degree at least $d+1$, then using Corollary 9, we conclude that $G\not\in\mathscr{C}_{k,d}$.
@@
There are at most $k^{2}$ branch sets containing a vertex of degree at least $d+1$, and therefore, at most $k^{2}\cdot(4k+7)^{2}$ branch sets that are at distance at most $32k+$ 3 to a vertex of degree at least $d+1$. As $h(k)$ is large enough, there is a $(k,d)$-safe branch set which, by Corollary 12, implies the existence of an irrelevant vertex. We iteratively eliminate irrelevant vertices until we are in one of Case 1 or Case 2.1.∎
We now do a quick complexity analysis of the algorithm. The algorithm of Theorem 5 runs in time $n^{O(1)}$. Then performing either Courcelle’s Theorem, finding $k^{2}$ branch set with vertices of degree $d+1$, or finding an irrelevant vertex can be done in linear time, i.e. $f(k,d)\cdot n$ for some computable function $n$. In the worst case, we might end up performing Case 2.2 up to $|G|$ many times before concluding via Case 1 or Case 2.1. Therefore, the overall complexity is $f(k,d)\cdot n^{c}$ for a computable function $f$ and constant $c$.
In the slightly more restrictive case where $G$ is planar, the algorithm of Theorem 6 runs in time $O(n^{2})$. this improves the complexity of the overall algorithm yielding a running time of $f(k,d)\cdot n^{3}$ for a computable function $f$.

### v4

@@ 3 The proof
Then, assume that $b$ is in $A_{\omega}\cup\bigcup\limits_{i\leq\omega}T_{i}$. Then $b$ is both $\leq_{G}$ and $\leq_{H}$-maximal.
Additionally, $S_{b}^{G}=S_{b}^{H}$ if $b$ is not adjacent to $a$, and $S_{b}^{G}=S_{b}^{H}\cup\{a\}$ otherwise. Note that as $a$ is $(k,d)$-safe, it has no neighbor of degree $d+1$, hence, if $b$ is adjacent to $a$ it has degree at most $d-1$ in $H$, and $|S_{b}^{H}|\leq d-1$ so $|S_{b}^{G}|\leq d$. For the last point of Definition 2.1, note that for any $v\in S^{G}_{b}$, $v$ also belongs to $A_{\omega}\cup\bigcup\limits_{i\leq\omega}T_{i}$, we then have $\{w~{}:~{}w<_{G}v\}=\{w~{}:~{}w<_{G}b\}=(m_{i})_{i<\omega}$.
Finally, we look at the case where $b$ is not in $A_{\omega}$, is not $m_{i}$ nor in $T_{i}$ for any $i\leq\omega$. In this case, we have that $b$ cannot have neighbors in $A_{\omega}$ nor any of the $T_{i}$ for $i\leq\omega$. To see this, assume that there is a $v\in T_{i}$ neighbor to $b$. This implies that $b$ is in $A_{i-1}$, as it is connected to $v$, the latter being in $A_{i-1}$, which is the connected component of $a$ in $G-\{m_{1},\ldots,m_{i-1}\}$. As $b\not\in T_{i}$, we have $m_{i}\leq_{H}b$ and $m_{i}\not\leq_{H}v$ which contradict that $\leq_{H}$ is an elimination order. This contradiction also holds if $b$ has a neighbor in $A_{\omega}$ as this would imply that $b\in A_{\omega}$.
@@
To conclude, we have that $\leq_{G}$ is indeed an elimination order to degree $d$ of depth $k$ for $G$. This ends the proof that $a$ is irrelevant.
We can now prove our main result, Theorem 1.1.
Proof 3.9
Let $k,d$ be two integers and $G$ be a connected $K_{5}$-minor free graph. We set $h(k):=(4k+7)\cdot k$. Let $g(\cdot)$ be the function from Theorem 2.4.
First run the algorithm of Theorem 2.4 with parameter $h(k)$ (or, if $G$ is planar, use the algorithm of Theorem 2.5).
@@
The algorithm outputs a tree-decomposition of width $g(h(k))$, or of width $5h(k)$ if $G$ is planar. We then use Courcelle’s Theorem (Theorem 2.6) to decide whether $G\in\mathscr{C}_{k,d}$.
@@
If more than $k^{2}$ branch sets contain a node of degree at least $d+1$, then using Corollary 3.3, we conclude that $G\not\in\mathscr{C}_{k,d}$.
@@
There are at most $k^{2}$ branch sets containing a vertex of degree at least $d+1$, and therefore, at most $k^{2}\cdot(4k+7)^{2}$ branch sets that are at distance at most $32k+$ 3 to a vertex of degree at least $d+1$. As $h(k)$ is large enough, there is a $(k,d)$-safe branch set which, by Corollary 3.7, implies the existence of an irrelevant vertex. We iteratively eliminate irrelevant vertices until we are in one of Case 1 or Case 2.1.
We now do a quick complexity analysis of the algorithm. The algorithm of Theorem 2.4 runs in time $n^{O(1)}$. Then performing either Courcelle’s Theorem, finding $k^{2}$ branch set with vertices of degree $d+1$, or finding an irrelevant vertex can be done in linear time, i.e. $f(k,d)\cdot n$ for some computable function $f$. In the worst case, we might end up performing Case 2.2 up to $|G|$ many times before concluding via Case 1 or Case 2.1. Therefore, the overall complexity is $f(k,d)\cdot n^{c}$ for a computable function $f$ and constant $c$.
In the slightly more restrictive case where $G$ is planar, the algorithm of Theorem 2.5 runs in time $O(n^{2})$. This improves the complexity of the overall algorithm yielding a total running time of $f(k,d)\cdot n^{3}$ for a computable function $f$.